OpenStudy (anonymous):

Given 2x-y=6, solve for y

OpenStudy (compassionate):

2x-y=6 x-y=3 y = 3 - x

OpenStudy (compassionate):

Now e have: -y = 3 - x y = -3 - x

OpenStudy (anonymous):

the answers that i didnt relize i put up are y=2x+6 y=2x-6 y=-2x+6 y=-2x-6